What Is The Slope Of A Line parallel to y = -2/5x + 4/5
Answer by nerdybill(7384)

What Is The Slope Of A Line parallel to y = -2/5x + 4/5
.
Two lines are parallel if it has the same slope.
because
y = -2/5x + 4/5
is in "slope-intercept" form:
y = mx + b
we see that the slope is -2/5
.
Answer: -2/5